OpenStudy (anonymous):

maximize the objective function 110x+170y, subject to the constraints. x+2y≤240 30x+20y≥240 x≤40 x≥0 y≥0

OpenStudy (anonymous):

you need to draw it on a piece of graph paper

OpenStudy (anonymous):

i need to find the maximum value of the function and the vlaue of x and y

OpenStudy (anonymous):

Are you sure 0 <= x <= 40?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

if you sketch the lines x +2y = 240, 30x + 20y = 240 and the vertical line x =0 and x = 40, and determine the region for x +2y <= 240, 30x + 20y >= 240, x<= 40, x>= 0, you will know your region of integration

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

do you have any calculus book with you?

OpenStudy (anonymous):

i do, but its not asking me to graph anything

OpenStudy (anonymous):

To find the local extremum of f in the interior of the region, you use second derivative test for function of two variables To get the extremum on the boundary of the region, you use Lagrange multiplier method